Caemello Anthony's chances of making a free throw on his first shot are 97%. If he makes his first shot then the chances of him makinf the second increases to 99%. But, i fhe misses his first shot, the chances of him making his second shot decreases to 96%. What is the probability that he makes both dhot? What is the probability that he makes exactly one shot?
1
Expert's answer
2011-05-19T07:16:17-0400
P(make 2 shots) = 0.97*0.99= 0.9603



P(make only one shot) = 0.97*(1-0.99) + (1-0.97)*0.96 = 0.0385